[figure: 9.7 cm]

On the 5d Band Polarization in R_{2}A_{17} (A=Fe, Co or Ni) Compounds

E. Burzo^{a}, P. Vlaic^{b} a Faculty of Physics, Babeş-Bolyai University, RO-3400 Cluj-Napoca, Romania

b Biophysics Department, University of Medicine and Pharmacy “ Iuliu Hatieganu” , 3400 Cluj-Napoca, Romania

Band structure calculations were performed on R_{2}Ni_{17}, R_{2}Co_{17} and R_{2}Fe_{17} compounds where R is a heavy rare-earth. The R(5d) polarizations are determined by 4f-5d local exchange as well as by 5d-3d short range exchange interactions with neighboring transition metal atoms. The first contribution is proportional to the De Gennes factor while the second one is determined by the number of A neighboring atoms as well as their magnetic contributions. The exchange interactions in the above systems are analyzed. The transition metal moments are discussed in correlation with their local environments.

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : poster

Corresponding author : E. Burzo

Address for correspondence : Faculty of Physics Babeş-Bolyai University RO-3400 Cluj-Napoca ROMANIA

Email address : burzo@phys.ubbcluj.ro